OpenStudy (anonymous):

Find the coordinates of the point (-5,7) when it is reflected across the line y=11.

OpenStudy (anonymous):

Use mid point formula.

OpenStudy (anonymous):

But don't you need more than one coordinate?

OpenStudy (anonymous):

mirror being mid point of the line joining the two points. \[\left\{ -5,11 \right\}\] now \[-5=\frac{x2-x1 }{ 2 }\] and \[11=\frac{ y2-y1 }{ 2 }\]

OpenStudy (anonymous):

you know x1 and y1 find x2 and y2

OpenStudy (anonymous):

Did you get it .

OpenStudy (anonymous):

I completely understand what you're saying, but I honestly have no idea how to solve for x2 and y2. I remember learning this, but I haven't used it in 2 years.

OpenStudy (anonymous):

did you get the idea behind mid point

OpenStudy (anonymous):

Yes I do!

OpenStudy (anonymous):

Then take the given point as the first point in the case i.e (-5,7)

OpenStudy (anonymous):

now use x1=-5 and y1=7

OpenStudy (anonymous):

now use the formula that I have already posted

OpenStudy (anonymous):

Wait, so wouldn't it be (-5,-7)?

OpenStudy (anonymous):

Nevermind, let me resolve quick.

OpenStudy (anonymous):

did you get is answer as (-5,15). If not I can give us an another approach